Hello, I have a question in regards to simplifying fractions.
Problem is 5/8x + 1/16x =11/16+x
I came up with x=88/80 and I was asked to simplify the answer, so because it is an improper fraction I turned it into a mixed number of 1 and(8/80), then I reduced it to 1 and(1/10). Can you please tell me if it is necessary to turn the problem into a mixed number?
Thanks in advance

Answer by Alan3354(69443)

It's not necessary unless someone asks for it.
11/10 is better than a mixed number, imo.
----------
5/8x + 1/16x =11/16+x
Is it (5/8)x ? Or 5/(8x) ?
---------------------------
11x/16 = x + (11/16)
11x = 16x + 11
-5x = 11
x = -11/5